A hot air balloon is filled with 1.21 x 10 L of an ideal gas on a cool morning (11 C). The air is heated to 115 C. What is
Answer #1

use relation

V1/T1= V2/T2

given initial volume V1 = 1.21 x10^6 L
initial temperature T1 = 11 C = 11+273 = 284 K
final temperature T1= 115 C = 388K
final volume V2 = ?

V2 = V1T2/T1 = 1.21 x10^6 L*388K/284 K = 1.653 x10^6 L

final volume =1.653 x10^6 L
*********
Thank you!!